OpenStudy (anonymous):

which equation is equivalent to 4 s+= t+2 s=t-2 s=4/t+2 s=t+2/4 or s=t+6

OpenStudy (anonymous):

help please

OpenStudy (mathmale):

something appears to be not quite right here: 4 s+= t+2 What do you mean by " + = " ?

OpenStudy (anonymous):

4s=t+2

OpenStudy (anonymous):

@mathmale

OpenStudy (kobeni-chan):

Given that the multiple choices have variable s by itself, you'll want to isolate s as you solve this. On the left hand side, it says '4s'. What operation do you need to make to get just s?

OpenStudy (anonymous):

\[4s=t+2\]divide both sides by 4

OpenStudy (anonymous):

\[\frac{ 4s }{ 4 }=\frac{ t+2 }{ 4 }\]

OpenStudy (anonymous):

\[s=?\]solve it @elocke